# A Scalable Design Approach to Resilient Architectures for Interconnected Cyber-Physical Systems: Safety Guarantees under Multiple Attacks

### v1

@@ I Introduction
Cyber-physical systems (CPSs) are exposed to malicious cyberattacks, as reported across numerous applications, including transportation [1] and power systems [2]. Such attacks may cause safety violations, with consequences ranging from equipment damage to significant risks to human safety. To this end, safety verification, safety-critical control design, and cyber-resilient architectures have been extensively studied [3, 4, 5, 6, 7, 8, 9, 10].
@@ III-A CI-Based Safety Guarantee
$\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)\hat{u}_{i}(x)\big)\;\geq\;\frac{c}{\tau},\quad\forall x\in\mathcal{C}\setminus\mathcal{C}_{c}$ (3)
@@
$\hat{\rho}_{i}=\inf_{x\in\mathcal{C}\setminus\mathcal{C}_{c},\,u_{i}\in\mathcal{U}_{i}}\left\{\frac{\partial h}{\partial x_{i}}\,g_{i}(x)(u_{i}-\hat{u}_{i})\right\}.$ (4)
The index $\hat{\rho}_{i}$ is a negative scalar that quantifies the worst-case rate of degradation of the safety function $h(x)$ resulting from deviations in the control input from the nominal policy in subsystem $\mathcal{S}_{i}$. A more negative value of $\hat{\rho}_{i}$ indicates a more critical subsystem, in the sense that its compromised input leads to faster safety degradation.
@@
$\begin{aligned} h(x(t))&=\int_{t_{0}}^{t}\dot{h}(x(t))\,dt+h(x_{0})\\ &\geq\int_{t_{0}}^{t_{f}}\Bigg(\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)~u_{i}\big)\Bigg)\,dt+c\\ &=\int_{t_{0}}^{t_{f}}\Bigg(\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)(u_{i}-\hat{u}_{i})+g_{i}(x)\hat{u}_{i}\big)\Bigg)\,dt+c\\ &=\int_{t_{0}}^{t_{f}}\Bigg(\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)~\hat{u}_{i}\big)\Bigg)\,dt\\ &\quad+\sum_{i=1}^{N}\Bigg(\int_{t_{0}}^{t_{f}}\frac{\partial h}{\partial x_{i}}g_{i}(x)(u_{i}-\hat{u}_{i})\,dt\Bigg)+c\end{aligned}$
First, note that the system must pass through the set $\mathcal{C}\setminus\mathcal{C}_{c}$ in order to become unsafe. By Assumption 1, we have $\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}(f_{i}(x)+g_{i}(x)~\hat{u}_{i})\geq 0$ for $x\in\mathcal{C}\setminus\mathcal{C}_{c}$. Furthermore, for each subsystem $S_{i}$, we have $u_{i}\neq\hat{u}_{i}$ only for $t(M(S_{i}))$ time-duration, as the maximum time an attacker can compromise a subsystem in a single attack cycle is bounded by its recovery time. Utilizing this fact as well as (5), we can write for each subsystem and for a single attack cycle that
$\int_{t_{0}}^{t_{f}}\frac{\partial h}{\partial x_{i}}g_{i}(x)({u}_{i}-\hat{u}_{i})~dt\geq\int_{0}^{t(M(S_{i}))}\rho_{i}~dt\geq\rho_{i}~t(M(S_{i}))$ (8)
@@
Therefore, if (6) is satisfied, then $h(x(t))\geq 0$, i.e. $x(t)\in\mathcal{C}$, which implies the system is safe.
@@ III-B Improvement of Bound via Segmentation
$\frac{\partial h}{\partial x_{i}}g_{i}(x)\left(u_{i}(x)-\hat{u}_{i}(x)\right)\geq\rho_{ij}~~~\forall x\in\mathcal{C}_{j},~\forall u_{i}\in\mathcal{U}_{i}$ (12)
@@
$\int_{t_{0}}^{t_{f}}\frac{\partial h}{\partial x_{i}}\,g_{i}(x)\,(u_{i}-\hat{u}_{i})\,dt\;\;\geq\;D_{i}(M(S_{i})).$ (15)
@@ IV Case Study
Figure 5: Time evolution of average temperature under partial overlapping attack with an overlap interval of $t_{\mathrm{ov}}=0.0018$ s. Subsystem $S_{1}$ is compromised after Phase 0 at $t=0.2$. During Phase 1, $S_{1}$ is recovered, and $S_{2}$ remains compromised. At Phase 2, $S_{2}$ is recovered and $S_{3}$ remains compromised. At Phase 3, $S_{3}$ is recovered. Finally, at Phase 4, the system returns to the set $\mathcal{C}_{c}$. Throughout the attack cycle, the system maintains safety.
@@ Acknowledgment
We thank the Network Security Lab (NSL) at the University of Washington. In particular, we thank Prof. Radha Poovendran and Prof. Luyao Niu (University of Washington), Prof. Andrew Clark (Washington University in St. Louis), Prof. Bhaskar Ramasubramanian (Western Washington University), and Prof. Sukarno Mertoguno (Georgia Institute of Technology) for their valuable suggestions and insights regarding this work.

### v2

@@ I Introduction
Cyber-physical systems (CPSs) are exposed to malicious cyberattacks, as reported across numerous applications, including transportation [8] and power systems [9]. Such attacks may cause safety violations, with consequences ranging from equipment damage to significant risks to human safety. To this end, safety verification, safety-critical control design, and cyber-resilient architectures have been extensively studied [2, 4, 3, 14, 6, 13, 22].
@@ III-A CI-Based Safety Guarantee
$\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)\hat{u}_{i}(x)\big)\;\geq\;\frac{c}{\tau},\quad~~~~\forall x\in\mathcal{C}\setminus\mathcal{C}_{c}$ (3)
@@
$\hat{\rho}_{i}=\inf_{x\in\mathcal{C}\setminus\mathcal{C}_{c},\,u_{i}\in\mathcal{U}_{i}}\left\{\frac{\partial h}{\partial x_{i}}\,g_{i}(x)(u_{i}-\hat{u}_{i})\right\}$ (4)
The index $\hat{\rho}_{i}$ is a negative scalar that quantifies the worst-case rate of degradation of the safety function $h(x)$ resulting from deviations in the control input from the nominal policy in subsystem $\mathcal{S}_{i}$. A more negative value of $\hat{\rho}_{i}$ indicates a more critical subsystem, in the sense that its compromised input leads to faster safety degradation. We emphasize that the criticality index is defined over the set $\mathcal{C}\setminus\mathcal{C}_{c}$, rather than the entire safety set $\mathcal{C}$, as this yields a less conservative result.
@@
$\begin{aligned} h(x(t))&=\int_{t_{0}}^{t}\dot{h}(x(t))\,dt+h(x(t_{0}))\\ &\geq\int_{t_{0}}^{t}\Bigg(\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)~u_{i}\big)\Bigg)\,dt+c\\ &=\int_{t_{0}}^{t}\Bigg(\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)(u_{i}-\hat{u}_{i})+g_{i}(x)\hat{u}_{i}\big)\Bigg)\,dt+c\\ &=\int_{t_{0}}^{t}\Bigg(\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}\big(f_{i}(x)+g_{i}(x)~\hat{u}_{i}\big)\Bigg)\,dt\\ &\quad+\sum_{i=1}^{N}\Bigg(\int_{t_{0}}^{t}\frac{\partial h}{\partial x_{i}}g_{i}(x)(u_{i}-\hat{u}_{i})\,dt\Bigg)+c\end{aligned}$
First, note that the system remains safe if it stays within the set $\mathcal{C}_{c}$ and it must pass through the set $\mathcal{C}\setminus\mathcal{C}_{c}$ in order to become unsafe. Therefore, to derive a sufficient condition for safety, we consider the worst-case scenario in which the system remains in the set $\{x\in\mathbb{R}^{n}:h(x)\leq c\}$ during the entire compromise period. Now, by Assumption 1, we have $\sum_{i=1}^{N}\frac{\partial h}{\partial x_{i}}(f_{i}(x)+g_{i}(x)~\hat{u}_{i})\geq 0$ for $x\in\mathcal{C}\setminus\mathcal{C}_{c}$. Furthermore, for each subsystem $S_{i}$, we have $u_{i}\neq\hat{u}_{i}$ only for $t(M(S_{i}))$ time-duration, as the maximum time an attacker can compromise a subsystem in a single attack cycle is bounded by its recovery time. Utilizing this fact together with (5), where $\rho_{i}$ is a negative scalar, we can write for each subsystem in a single attack cycle that
$\int_{t_{0}}^{t}\frac{\partial h}{\partial x_{i}}g_{i}(x)({u}_{i}-\hat{u}_{i})~dt\geq\int_{0}^{t(M(S_{i}))}\rho_{i}~dt\geq\rho_{i}~t(M(S_{i}))$ (8)
@@
Therefore, if (6) is satisfied, then $h(x(t))\geq 0$, i.e. $x(t)\in\mathcal{C}$, which implies the system is safe for all $t\in[t_{0},t_{f}]$.
@@ III-B Improvement of Bound via Segmentation
$\frac{\partial h}{\partial x_{i}}g_{i}(x)\left(u_{i}(x)-\hat{u}_{i}(x)\right)\geq\rho_{ij},~~~\forall x\in\mathcal{C}_{j},~\forall u_{i}\in\mathcal{U}_{i}$ (12)
@@
$\int_{t_{0}}^{t}\frac{\partial h}{\partial x_{i}}\,g_{i}(x)\,(u_{i}-\hat{u}_{i})\,dt\;\;\geq\;D_{i}(M(S_{i}))$ (15)
@@ IV Case Study
Figure 5: Time evolution of average temperature under partial overlapping attack with an overlap interval of $t_{\mathrm{ov}}=0.0018$ s. Subsystem $S_{1}$ is compromised after Phase 0 at $t=0.2$. During Phase 1, $S_{1}$ is recovered, and $S_{2}$ remains compromised. At Phase 2, $S_{2}$ is recovered and $S_{3}$ remains compromised. At Phase 3, $S_{3}$ is recovered. At Phase 4, the system returns to the set $\mathcal{C}_{c}$. Throughout the attack cycle, the system maintains safety.
@@ Acknowledgment
We thank the Network Security Lab (NSL) at the University of Washington. In particular, we thank Prof. Radha Poovendran and Prof. Luyao Niu (University of Washington), Prof. Andrew Clark (Washington University in St. Louis), Prof. Bhaskar Ramasubramanian (Western Washington University), and Prof. Sukarno Mertoguno (Georgia Institute of Technology) for their valuable suggestions on this work.